OpenStudy (anonymous):

Choose the equation of the line passing through the point (-3, 1) and parallel to y = -x - 4.

OpenStudy (anonymous):

Firstly find the slope of required line.. This you can do by just finding the slope of given line equation: ie y = -x + 4 Can you find the slope of y = -x + 4 ?

OpenStudy (anonymous):

Remember : Slope of parallel lines are equal..

OpenStudy (anonymous):

oh okay, i dont really know how to do that yet im aking th epretest now i havent been taught the lesson

OpenStudy (anonymous):

So, you don't know how to find slope of an equation?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

do you know how to solve this?